# Learnable In-Context Vector for Visual Question Answering

### v1

@@ Abstract
As language models continue to scale, Large Language Models (LLMs) have exhibited emerging capabilities in In-Context Learning (ICL), enabling them to solve language tasks by prefixing a few in-context demonstrations (ICDs) as context. Inspired by these advancements, researchers have extended these techniques to develop Large Multimodal Models (LMMs) with ICL capabilities. However, applying ICL usually faces two major challenges: 1) using more ICDs will largely increase the inference time and 2) the performance is sensitive to the selection of ICDs. These challenges are further exacerbated in LMMs due to the integration of multiple data types and the combinational complexity of multimodal ICDs. Recently, to address these challenges, some NLP studies introduce non-learnable In-Context Vectors (ICVs) which extract useful task information from ICDs into a single vector and then insert it into the LLM to help solve the corresponding task. However, although useful in simple NLP tasks, these non-learnable methods fail to handle complex multimodal tasks like Visual Question Answering (VQA). In this study, we propose Learnable ICV (L-ICV) to distill essential task information from demonstrations, improving ICL performance in LMMs. Experiments show that L-ICV can significantly reduce computational costs while enhancing accuracy in VQA tasks compared to traditional ICL and other non-learnable ICV methods.
@@ 1 Introduction
As language models continue to scale up, Large Language Models (LLMs) [1, 2, 3] have demonstrated emerging capabilities in In-Context Learning (ICL) [4]: these models can solve language tasks when provided with a few similar examples, termed in-context demonstrations (ICDs), as context. Unlike traditional task-specific fine-tuning, ICL achieves comparable performance without necessitating updates to millions or trillions of model parameters [5]. By prefixing just a handful of data samples to the query input, ICL configures a model’s behavior to produce the corresponding output, thus facilitating rapid adaptation across a wide range of downstream tasks. Inspired by these advancements in the language domain, researchers have extended these techniques to develop Large Multimodal Models (LMMs) with ICL capabilities [6, 7, 8].
@@
Figure 1: (a) Conventional ICL is more sensitive to the ICD selection and requires more inference time. (b) L-ICV is more robust and reduces inference time by inputting a shift vector.
@@
In this study, to make ICVs abstract more useful VQA task information, we try to distill the task information implied in demonstrations into a single Learnable ICV (L-ICV). Our method is motivated by the observation [18] that ICL can be treated as a process of “shifting" the direction of the latent states of query towards the target, i.e., adding this latent state with a shift vector. Then we hope to learn suitable ICVs to replace the ICDs during inference to shift the direction. To achieve this, we train this L-ICV by minimizing the output distributions of a LMM got by only using this L-ICV and by inputting a few demonstrations. During training, we use different 32-shot randomly sampled demonstrations for different queries to distill task knowledge. Then L-ICV is encouraged to capture the most essential task information from these different combinations by removing the individual characteristics of demonstrations. Moreover,  [22] finds that during ICL, different layers of LLM have diverse roles in addressing demonstrations. Then in our method, each layer is assigned with a unique ICV to capture more fine-grained task information.
Our L-ICV inherits the efficiency of previous non-learnable ICVs, i.e., during inference, under the same performance conditions, L-ICV only needs 1/24.97 FLOPs number of 32-shot ICL. Additionally, in VQAV2/OKVQA, L-ICV improves accuracy by 2.36/1.6 compared to 32-shot ICL. We also compare L-ICV to LoRA [23] that when comparable trainable parameters are used, L-ICV requires much fewer training samples than LoRA (500 vs. 8000) to achieve satisfactory performance. Besides, we design lots of analytical experiments to validate whether L-ICV can better shift the hidden states of queries to the target direction and analyze why previous non-learnable methods fail to solve VQA.
@@ 2 Related Work
ICL in LLM: Prompt engineering allows LLMs to tackle specific tasks without requiring fine-tuning [24, 25, 26]. A specific form of this approach, ICL, further improves these capabilities by creating prompts that include several demonstrations. ICL has already demonstrated superior performance and good generalization on many tasks [27, 5, 28], and can be easily adapted to downstream tasks. However, the use of ICL faces several issues: first, ICL is very sensitive to the selection and arrangement order of demonstrations [4, 9, 10, 11, 29]; poor demonstrations can severely impact ICL performance. Second, too many demonstrations can significantly slow down the inference speed of LLMs [30, 31]. While ICVs can effectively address these two issues, as it can use only queries as input to the model while preserving ICL performance, without the need for demonstrations as input.
@@
ICL in LMM: As the performance of LLMs continues to improve, an increasing number of researchers begin to adapt LLMs to the multimodal domain [32, 33, 34, 35, 36]. Relying on the powerful inference capabilities of LLMs, some LMMs have started to exhibit ICL capabilities, such as Flamingo [6] and IDEFICS [7]. Moreover, these models have significantly enhanced their ICL capabilities by concatenating multiple samples as contextual information during the training process. Currently, researchers mainly focus on how to configure demonstrations to address the sensitivity of ICL performance in LMMs. [16, 17] have respectively adopted heuristic retrieval methods for selecting demonstrations in Image Captioning and VQA. However, no researchers have yet extracted ICV from LMMs and evalutaed it. Therefore, the effectiveness of ICV in LMMs still needs further exploration. Considering that the IDEFICS model shares the same model structure as Flamingo and possesses stronger ICL capabilities, we primarily focus on valid our method on the IDEFICS model.
@@
Figure 2: The L-ICV training pipeline: (a) The distribution $\mathcal{P}(\bm{\hat{x}}|\bm{V},\bm{\alpha};\mathcal{M})$ of LMMs output when using L-ICV. (b) Adding L-ICV into the representations of the query to simulate the shift effect brought by demonstrations. (c) The distribution $\mathcal{P}(\bm{\hat{x}}|\bm{X}_{D};\mathcal{M})$ of LMMs output when using demonstrations.
3 Learnable In-Context Vector
@@
Here we show how to derive the formulation of the shift vector from Self-Attention (SA) mechanism and then introduce how to design L-ICV based on this formulation. Generally, to implement In-Context Learning (ICL) using a language or multimodal model (LLM/LMM) $\mathcal{M}$, the input has the following form: $\bm{X}=\{\bm{X}_{D},\hat{\bm{x}}\}$, where $\bm{X}_{D}=\{\bm{x}_{1},...,\bm{x}_{k}\}$ represents the concatenation of $k$ In-Context Demonstrations (ICDs), and $\hat{\bm{x}}$ denotes the query input, as shown in Figure 2. Given $\bm{X}$ as Key and Value, for each token $\hat{x}_{i}$ of $\hat{\bm{x}}$, applying Self-Attention (SA) once yields:
@@
$=\frac{Z_{1}}{Z_{1}+Z_{2}}\frac{\exp(\hat{x}_{i}\bm{X}_{D}^{\top})\bm{X}_{D}}{Z1}+\frac{Z_{2}}{Z_{1}+Z_{2}}\frac{\exp(\hat{x}_{i}\hat{\bm{x}}^{\top})}{Z_{2}}$
@@
We propose a novel method that involves a Learnable In-Context Vector (L-ICV) to simulate the ICL process without actual demonstrations. This approach aims to abstract general task information from demonstrations, enabling it to shift the model’s representation toward the direction influenced by the ICDs. Figure 2 shows the training pipeline of L-ICV. The L-ICV training dataset, denoted as $\mathcal{D}=\{\bm{d}_{1},\dots,\bm{d}_{N}\}$, is a subset of the VQA dataset training split, created by randomly selecting $N$ question-answer pairs from it. We use each training sample $\bm{d}_{i}$ to simulate the query sample $\bm{\hat{x}}$ in ICL, and randomly select $k$ demonstrations from $\mathcal{D}\setminus\{\bm{d}_{i}\}$ for it. Additionally,  [22] shows that during ICL, each layer of an LLM performs a distinct role. Motivated by this, we assume that for LMM, each layer also requires a specific shift direction. We assign a learnable vector $\bm{v}_{l}$ and a weight factor $\alpha_{l}$ for each layer $l$ to learn the unique shift effect. Our final L-ICV comprises of the vector set $\bm{V}$ and the corresponding weight factor set $\bm{\alpha}$ as:
@@
where $L$ is the number of layers. To train $\bm{V}$ and $\bm{\alpha}$, we align the distribution of the model’s outputs for the query when shifted by demonstrations, $\mathcal{P}(\bm{\hat{x}}|\bm{X}_{D};\mathcal{M})$, with that shifted by our L-ICV, $\mathcal{P}(\bm{\hat{x}}|\bm{V},\bm{\alpha};\mathcal{M})$. This alignment is achieved by minimizing the Kullback-Leibler (KL) divergence:
@@
To obtain the output of $\bm{\hat{x}}$ by using L-ICV, at the $l$-th layer, we use the vector $\bm{v}_{l}$ to shift original representation and get: $h(\hat{x}_{i})^{\prime}=h(\hat{x}_{i})+\alpha_{l}\bm{v}_{l}$, which is shown in Figure 2(a). After applying L-ICV to shift the representations at each layer, we obtain the output distribution $\mathcal{P}(\bm{\hat{x}}|\bm{V},\bm{\alpha};\mathcal{M})$. Notably, during training, $\bm{X}_{D}$ for each query $\hat{\bm{x}}$ includes randomly sampled 32-shot demonstrations. This strategy encourages our L-ICV to extract the most useful common information from various demonstration combinations and prevents it from being influenced by the individual characteristics of certain demonstrations.
In addition, to facilitate the L-ICV in acquiring more task-specific information, we also optimize the $\mathcal{P}(\bm{\hat{x}}|\bm{V},\bm{\alpha};\mathcal{M})$ with the ground truth by $\mathcal{L}_{\text{gt }}$. Thus, the overall loss $\mathcal{L}$ is defined as:
@@ 4.1 Setting and implementation details
Model and Dataset: We evaluate our approach using the IDEFICS-9B model [7] across two datasets: VQAv2 [37] and OKVQA [38]. VQAv2 emphasizes open-ended VQA tasks, encompassing $4,437,570$ question-answer pairs in its training split, supplemented by an additional $2,143,540$ pairs in the validation split. OKVQA is a large-scale dataset designed for models that require external knowledge to answer questions. It consists of $14,055$ question-answer pairs, with $9,009$ allocated for training and $5,046$ for validation. For both VQAv2 and OKVQA datasets, We train our L-ICV on $8,000$ pairs from each training set. Due to computational resource limitations, we randomly sample $10,000$ question-answer pairs from the VQAv2 validation split for evaluation [16]. For OKVQA, we utilize the entire validation split.
L-ICV Setting: During training, we assign 32-shot demonstrations for each query, enabling L-ICV to acquire better directions of shifting vectors for VQA tasks. The $\bm{v}_{i}$ is initialized using a normal distribution with a mean of 00 and a standard deviation of $0.01$, and all $\alpha_{i}$ are initialized to $0.1$. More detailed training parameters can be found in Appendix.
@@ 4.2.1 Compared Methods
LoRA [23]: This method finetunes the LMMs with the same number of samples of training L-ICV. We add the LoRA module in the token classification head of the last layer. In this way, the number of trainable parameters is comparable to that of L-ICV.
Table 1: Accuracy (%) with Different ICVs Methods and Finetuning Methods, where numbers in parentheses indicate multiples of L-ICV trainable parameters.
Zero-Shot 32-shot ICL TV FV PCA-ICV LoRA L-ICV (Ours)
@@
We present performance comparisons with various methods in Table 1. Certain existing methods show only marginal improvements over Zero-Shot baselines, e.g., FV improves by 0.96/0.48 on VQAv2/OKVQA and PCA-ICV improves by 0.04 on OKVQA. Besides, we observe that all the previous non-learnable ICV methods do not reach the performance of the standard 32-shot ICL, e.g., the best non-learnable method, TV, is still 12.5/15.8 lower than 32-shot ICL on VQAv2/OKVQA. In contrast, our L-ICV achieves an accuracy improvement of 2.36 on VQAv2 and 1.6 on OKVQA over 32-shot ICL. These results highlight the inefficacy of non-learnable methods in capturing essential task-specific information for VQA, whereas L-ICV, by leveraging diverse 32-shot ICL demonstrations for each query during training, manages to abstract useful task information effectively. We further show that our L-ICV outperforms LoRA with less trainable parameters, suggesting L-ICV can abstract task information more efficiently.
Figure 3 displays the efficiency of L-ICV during inference compared to other methods. We average the FLOPs and actual inference time consumption per forward pass over 1000 randomly sampled queries. [^1] For detailed hardware information, refer to Appendix. We observe that L-ICV only needs 1/24.97 FLOPs and 1/8.25 inference time of 32-shot ICL per forward pass. Additionally, L-ICV maintains almost the same inference speed as Zero-Shot. These comparisons validate the efficiency of L-ICV during inference.
Figure 3: The total number of FLOPs and real inference time consumption of ICL, Zero-Shot, L-ICV for 1000 query samples.

### v3

# LIVE: Learnable In-Context Vector for Visual Question Answering

@@ Abstract
As language models continue to scale, Large Language Models (LLMs) have exhibited emerging capabilities in In-Context Learning (ICL), enabling them to solve language tasks by prefixing a few in-context demonstrations (ICDs) as context. Inspired by these advancements, researchers have extended these techniques to develop Large Multimodal Models (LMMs) with ICL capabilities. However, applying ICL usually faces two major challenges: 1) using more ICDs will largely increase the inference time and 2) the performance is sensitive to the selection of ICDs. These challenges are further exacerbated in LMMs due to the integration of multiple data types and the combinational complexity of multimodal ICDs. Recently, to address these challenges, some NLP studies introduce non-learnable In-Context Vectors (ICVs) which extract useful task information from ICDs into a single vector and then insert it into the LLM to help solve the corresponding task. However, although useful in simple NLP tasks, these non-learnable methods fail to handle complex multimodal tasks like Visual Question Answering (VQA). In this study, we propose Learnable In-Context Vector (LIVE) to distill essential task information from demonstrations, improving ICL performance in LMMs. Experiments show that LIVE can significantly reduce computational costs while enhancing accuracy in VQA tasks compared to traditional ICL and other non-learnable ICV methods. The code is available at https://github.com/ForJadeForest/LIVE-Learnable-In-Context-Vector.
@@ 1 Introduction
As language models continue to scale up, Large Language Models (LLMs) [1, 2, 3] have demonstrated emerging capabilities in In-Context Learning (ICL) [4]: these models can solve language tasks when provided with a few similar examples, termed in-context demonstrations (ICDs), as context. Unlike traditional task-specific fine-tuning, ICL, a efficient method to adapt LLM to downstream task [5, 6], achieves comparable performance without necessitating updates to millions or trillions of model parameters [7]. By prefixing just a handful of data samples to the query input, ICL configures a model’s behavior to produce the corresponding output, thus facilitating rapid adaptation across a wide range of downstream tasks. Inspired by these advancements in the language domain, researchers have extended these techniques to develop Large Multimodal Models (LMMs) with ICL capabilities [8, 9, 10].
@@
Figure 1: (a) Conventional ICL is more sensitive to the ICD selection and requires more inference time. (b) LIVE is more robust and reduces inference time by inputting a shift vector.
@@
In this study, to make ICVs abstract more useful VQA task information, we try to distill the task information implied in demonstrations into a single Learnable In-Context Vector (LIVE). Our method is motivated by the observation [20] that ICL can be treated as a process of “shifting" the direction of the latent states of query towards the target, i.e., adding this latent state with a shift vector. Then we hope to learn suitable ICVs to replace the ICDs during inference to shift the direction. To achieve this, we train LIVE by minimizing the output distributions of a LMM got by only using LIVE and by inputting a few demonstrations. During training, we use different 32-shot randomly sampled demonstrations for different queries to distill task knowledge. Then LIVE is encouraged to capture the most essential task information from these different combinations by removing the individual characteristics of demonstrations. Moreover,  [24] finds that during ICL, different layers of LLM have diverse roles in addressing demonstrations. Then in our method, each layer is assigned with a unique ICV to capture more fine-grained task information.
Our LIVE inherits the efficiency of previous non-learnable ICVs, i.e., during inference, under the same performance conditions, LIVE only needs 1/24.97 FLOPs number of 32-shot ICL. Additionally, in VQAV2/OKVQA, LIVE improves accuracy by 2.36/1.6 compared to 32-shot ICL. We also compare LIVE to LoRA [25] that when comparable trainable parameters are used, LIVE requires much fewer training samples than LoRA (500 vs. 8000) to achieve satisfactory performance. Besides, we design lots of analytical experiments to validate whether LIVE can better shift the hidden states of queries to the target direction and analyze why previous non-learnable methods fail to solve VQA.
@@ 2 Related Work
ICL in LLM: Prompt engineering allows LLMs to tackle specific tasks without requiring fine-tuning [26, 27, 28, 29, 30, 31, 32, 33]. A specific form of this approach, ICL, further improves these capabilities by creating prompts that include several demonstrations. ICL has already demonstrated superior performance and good generalization on many tasks [34, 7, 35, 36], and can be easily adapted to downstream tasks. However, the use of ICL faces several issues: first, ICL is very sensitive to the selection and arrangement order of demonstrations [4, 11, 12, 13, 37, 38, 39]; poor demonstrations can severely impact ICL performance. Second, too many demonstrations can significantly slow down the inference speed of LLMs [40]. While ICVs can effectively address these two issues, as it can use only queries as input to the model while preserving ICL performance, without the need for demonstrations as input.
@@
ICL in LMM: As the performance of LLMs continues to improve, an increasing number of researchers begin to adapt LLMs to the multimodal domain [41, 42, 43, 44, 45, 46]. Relying on the powerful inference capabilities of LLMs, some LMMs have started to exhibit ICL capabilities, such as Flamingo [8] and IDEFICS [9]. Moreover, these models have significantly enhanced their ICL capabilities by concatenating multiple samples as contextual information during the training process. Currently, researchers mainly focus on how to configure demonstrations to address the sensitivity of ICL performance in LMMs. [18, 19] have respectively adopted heuristic retrieval methods for selecting demonstrations in Image Captioning and VQA. However, no researchers have yet extracted ICV from LMMs and evalutaed it. Therefore, the effectiveness of ICV in LMMs still needs further exploration. Considering that the IDEFICS model shares the same model structure as Flamingo and possesses stronger ICL capabilities, we primarily focus on valid our method on the IDEFICS model.
@@
Figure 2: The LIVE training pipeline: (a) The distribution $\mathcal{P}(\bm{\hat{x}}|\bm{V},\bm{\alpha};\mathcal{M})$ of LMMs output when using LIVE. (b) Adding LIVE into the representations of the query to simulate the shift effect brought by demonstrations. (c) The distribution $\mathcal{P}(\bm{\hat{x}}|\bm{X}_{D};\mathcal{M})$ of LMMs output when using demonstrations.
3 LIVE: Learnable In-Context Vector
@@
Here we show how to derive the formulation of the shift vector from Self-Attention (SA) mechanism and then introduce how to design LIVE based on this formulation. Generally, to implement In-Context Learning (ICL) using a language or multimodal model (LLM/LMM) $\mathcal{M}$, the input has the following form: $\bm{X}=\{\bm{X}_{D},\hat{\bm{x}}\}$, where $\bm{X}_{D}=\{\bm{x}_{1},...,\bm{x}_{k}\}$ represents the concatenation of $k$ In-Context Demonstrations (ICDs), and $\hat{\bm{x}}$ denotes the query input, as shown in Figure 2. Given $\bm{X}$ as Key and Value, for each token $\hat{x}_{i}$ of $\hat{\bm{x}}$, applying Self-Attention (SA) once yields:
@@
$=\frac{Z_{1}}{Z_{1}+Z_{2}}\frac{\exp(\hat{x}_{i}\bm{X}_{D}^{\top})\bm{X}_{D}}{Z_{1}}+\frac{Z_{2}}{Z_{1}+Z_{2}}\frac{\exp(\hat{x}_{i}\hat{\bm{x}}^{\top})}{Z_{2}}$
@@
We propose a novel method that involves a Learnable In-Context Vector (LIVE) to simulate the ICL process without actual demonstrations. This approach aims to abstract general task information from demonstrations, enabling it to shift the model’s representation toward the direction influenced by the ICDs. Figure 2 shows the training pipeline of LIVE. The LIVE training dataset, denoted as $\mathcal{D}=\{\bm{d}_{1},\dots,\bm{d}_{N}\}$, is a subset of the VQA dataset training split, created by randomly selecting $N$ question-answer pairs from it. We use each training sample $\bm{d}_{i}$ to simulate the query sample $\bm{\hat{x}}$ in ICL, and randomly select $k$ demonstrations from $\mathcal{D}\setminus\{\bm{d}_{i}\}$ for it. Additionally,  [24] shows that during ICL, each layer of an LLM performs a distinct role. Motivated by this, we assume that for LMM, each layer also requires a specific shift direction. We assign a learnable vector $\bm{v}_{l}$ and a weight factor $\alpha_{l}$ for each layer $l$ to learn the unique shift effect. Our final LIVE comprises of the vector set $\bm{V}$ and the corresponding weight factor set $\bm{\alpha}$ as:
@@
where $L$ is the number of layers. To train $\bm{V}$ and $\bm{\alpha}$, we align the distribution of the model’s outputs for the query when shifted by demonstrations, $\mathcal{P}(\bm{\hat{x}}|\bm{X}_{D};\mathcal{M})$, with that shifted by our LIVE, $\mathcal{P}(\bm{\hat{x}}|\bm{V},\bm{\alpha};\mathcal{M})$. This alignment is achieved by minimizing the Kullback-Leibler (KL) divergence:
@@
To obtain the output of $\bm{\hat{x}}$ by using LIVE, we follow [20, 22], we use the vector $\bm{v}_{l}$ to shift the each layer’s output representation $\mathcal{M}_{l}(\hat{x}_{i})$ and get: $\mathcal{M}_{l}(\hat{x}_{i})^{\prime}=\mathcal{M}_{l}(\hat{x}_{i})+\alpha_{l}\bm{v}_{l}$, which is shown in Figure 2(a). After applying LIVE to shift the representations at each layer, we obtain the output distribution $\mathcal{P}(\bm{\hat{x}}|\bm{V},\bm{\alpha};\mathcal{M})$. Notably, during training, $\bm{X}_{D}$ for each query $\hat{\bm{x}}$ includes randomly sampled 32-shot demonstrations. This strategy encourages our LIVE to extract the most useful common information from various demonstration combinations and prevents it from being influenced by the individual characteristics of certain demonstrations.
In addition, to facilitate the LIVE in acquiring more task-specific information, we also optimize the $\mathcal{P}(\bm{\hat{x}}|\bm{V},\bm{\alpha};\mathcal{M})$ with the ground truth by $\mathcal{L}_{\text{gt }}$. Thus, the overall loss $\mathcal{L}$ is defined as:
@@ 4.1 Setting and implementation details
Model and Dataset: We evaluate our approach using the IDEFICS-9B model [9] across two datasets: VQAv2 [47] and OKVQA [48]. VQAv2 emphasizes open-ended VQA tasks, encompassing $4,437,570$ question-answer pairs in its training split, supplemented by an additional $2,143,540$ pairs in the validation split. OKVQA is a large-scale dataset designed for models that require external knowledge to answer questions. It consists of $14,055$ question-answer pairs, with $9,009$ allocated for training and $5,046$ for validation. For both VQAv2 and OKVQA datasets, We train our LIVE on $8,000$ pairs from each training set. Due to computational resource limitations, we randomly sample $10,000$ question-answer pairs from the VQAv2 validation split for evaluation [18]. For OKVQA, we utilize the entire validation split.
LIVE Setting: During training, we assign 32-shot demonstrations for each query, enabling LIVE to acquire better directions of shifting vectors for VQA tasks. The $\bm{v}_{i}$ is initialized using a normal distribution with a mean of 00 and a standard deviation of $0.01$, and all $\alpha_{i}$ are initialized to $0.1$. More detailed training parameters can be found in Appendix.
@@ 4.2.1 Compared Methods
LoRA [25]: This method finetunes the LMMs with the same number of samples of training LIVE. We add the LoRA module in the token classification head of the last layer. In this way, the number of trainable parameters is comparable to that of LIVE.
Table 1: Accuracy (%) with Different ICVs Methods and Finetuning Methods, where numbers in parentheses indicate multiples of LIVE trainable parameters.
Zero-Shot 32-shot ICL TV FV PCA-ICV LoRA LIVE (Ours)
@@
We present performance comparisons with various methods in Table 1. Certain existing methods show only marginal improvements over Zero-Shot baselines, e.g., FV improves by 0.96/0.48 on VQAv2/OKVQA and PCA-ICV improves by 0.04 on OKVQA. Besides, we observe that all the previous non-learnable ICV methods do not reach the performance of the standard 32-shot ICL, e.g., the best non-learnable method, TV, is still 12.5/15.8 lower than 32-shot ICL on VQAv2/OKVQA. In contrast, our LIVE achieves an accuracy improvement of 2.36 on VQAv2 and 1.6 on OKVQA over 32-shot ICL. These results highlight the inefficacy of non-learnable methods in capturing essential task-specific information for VQA, whereas LIVE, by leveraging diverse 32-shot ICL demonstrations for each query during training, manages to abstract useful task information effectively. We further show that our LIVE outperforms LoRA with less trainable parameters, suggesting LIVE can abstract task information more efficiently.
Figure 3 displays the efficiency of LIVE during inference compared to other methods. We average the FLOPs and actual inference time consumption per forward pass over 1000 randomly sampled queries. [^1] For detailed hardware information, refer to Appendix. We observe that LIVE only needs 1/24.97 FLOPs and 1/8.25 inference time of 32-shot ICL per forward pass. Additionally, LIVE maintains almost the same inference speed as Zero-Shot. These comparisons validate the efficiency of LIVE during inference.
Figure 3: The total number of FLOPs and real inference time consumption of ICL, Zero-Shot, LIVE for 1000 query samples.